A standard cast from pipe inner diameter = 50mm and outer diameter is insulated with 85% magnesium insulation (k = 0.02w/mc) temperature at interface between the pipe and insulation is 300 degree celcius, the allowable heat loss through pipe is 600w/m length of pipe


1
Expert's answer
2022-03-23T01:37:06-0400

 Heat loss :

"\frac{Q}{t}=\frac{kA(T_2-T1)}{d}=\frac{0.02\times (300)}{50} = 0.12 \frac{w}{m^2}"